According to a 2009 Reader's Digest article, people throw away approximately 11% of what they buy at the grocery store. Assume this is the true proportion and you plan to randomly survey 75 grocery shoppers to investigate their behavior. What is the probability that the sample proportion exceeds 0.12?

Note: You should carefully round any z-values you calculate to 4 decimal places to match wamap's approach and calculations.
Answer #1
for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.110
sample size       =n= 75
std error of proportion=σp=√(p*(1-p)/n)= 0.0361

probability that the sample proportion exceeds 0.12 :

probability = P(X>0.12) = P(Z>0.28)= 1-P(Z<0.28)= 1-0.6103= 0.3897